A buffer solution is prepared by dissolving 1.50gram of benzoic acid HC7H5O2 and sodium benzoate NaC7H5O2 in 150 mL of Solution a. what is the PH of this buffer Solution? ( ka =6.3×10^-5)
1
Expert's answer
2021-03-10T06:54:15-0500

mass of Benzoic acid = 1.50g

number of moles of Benzoic acid = 1.50/122.12 = 0.0123 moles

molar concentration [HA] = 0.0123/150 × 1000 = 0.082 M


Ka = 6.3 × 10-5

pKa = 4.2


from,

"pK_a = pH - \log( \dfrac{A^-}{HA})"


"pH = pK_a + \log( \dfrac{A^-}{HA})"


"pH =4.2 + \log(\dfrac{A^-}{0.082})"